OpenStudy (anonymous):

How can x2 + 3x + 1 = 2x2 + 2x + 3 be set up as a system of equations?

OpenStudy (anonymous):

let y = left sign and y = right side

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

y = x2 + 3x + 1 and y = 2x2 + 2x + 3

OpenStudy (phi):

I would just simplify them into one equation = 0 then factor it.

OpenStudy (anonymous):

wpuld it be y = x2 + 3x + 1 y = 2x2 + 2x + 3

OpenStudy (anonymous):

@phi

OpenStudy (phi):

if you want a system of equations, you know they both equal the same number, call it c: x^2 + 3x + 1=c 2x^2 + 2x + 3=c it that what you want to do? or do you want to solve it?

OpenStudy (anonymous):

yes i just need to sepearte the current equations into a system i don't need to solve

OpenStudy (owlcoffee):

I'll explain, and let sourwing do the rest. If we look back in the definition of "system of equation", it is a set of two or more equations with various variables wich conform a mathematical problem that consists in fining the value/values that satisfies both equations. so let's say I have a generalized case, where I have "m" equations with "n" variables, it's generalized form is: \[1) F _{1} (x _{1}...x _{n})=0\] \[2)F _{2}(x _{1}...x _{n})=0\] . . . \[m)F _{m}(x _{1}...x _{n}) = 0\] where F1,F2...Fm are functions of the variables, the solution will be all the Fi that has values wich verifies the equality. Saying so, all the F's must be equal. So say I have an equality: \[Ax ^{2}+Bx+C=A'x ^{2}+B'x+C'\] by the definition I gave youn it'll be totally fine since both equations are equal, and C, C' are constant, to write them as: \[1)Ax ^{2}+Bx=-C\] \[2)A'x ^{2}+B'x=-C'\]

OpenStudy (anonymous):

thank you!!

OpenStudy (owlcoffee):

System of equations takes a very deep understanding as you continue learning math, and is a very useful tool, be sure to master them as quicky as possible ;)